(3D ^ 2 + D ^ 2 + D) * Z = X + 5Y

1
Expert's answer
2021-02-24T06:41:46-0500

3D2+D2+D=D(4D+1)=0

Roots: 0,-1/4

CF: c1e0+c2e-1/4=c1+c2e-1/4

PI: "\frac{x+5y}{4D+1}=\frac{x+5y}{4(D+1/4)}(1)"

"\frac{x+5y}{4} \frac{1}{D} {1/4}=\frac{x+5y}{16}xy"

"z=c_1+c_2e^{-1/4}+\frac{x+5y}{16}xy"